In a college of 300 students , every student reads 5 newspapers and every newspaper is read by 60 students . The number of newspapers is ? [ IIT 1998] .
The solution is given as , let x be the no.of newspapers therefore 60*x = 300 * 5 . This question is given in topic sets . Pls explain how it is equated and why . How 60x=300*5 .

Grade:12th Pass

8 Answers

Sher Mohammad IIT Delhi
10 years ago
total different newspapers possible 300 x 5 = 1500 newspapers.
total different newspapers if a newspaper is read by 60 student, 1500 / 60 = 25

Sher Mohammad
B.Tech, IIT Delhi
SANIKA
10 years ago
Pls explain further on what basis it is equated as 300 x 5 = 60x
Rakesh Siraswal
7 years ago
Total no. of  students= 300.
Let total no. of news papers= n.
Given every student reads 5 news papers  and every newspaper is read by 60 students.
Total no. of  newspaper reading by every student = 300*5 = 1500 .
Total no. of newspaper reading which each paper will get = 60*n .
which implies that  60*n = 300 *5
                   =>    60 *n =1500
                     =>   n = 25.
Total no. of newspapers is 25.
Pratyush Ranjan
7 years ago
one student reads 5 newspaper (we can say 5 types of newspaper)
300 students will also read 5types of newspaper
one newspaper is read by 60 students
so 300 students will read 5newspaper of one type
so 5newspaper of 5types =25 
abhisek
7 years ago
There are 300 students... so if we divide them in a section of 60 student...there will be 5 groups.. if there will be 5 news paper so..every group will read one news paper and then interchange.. so like wise there will be at least 5 news paper needed
Govind murali
6 years ago
I think i
Krishna
5 years ago
Let total newspapers be n.
Now every student reads 5 types of newspapers (N1,N2,N3,N4,N5)
And each newspaper is read by 60 students.
Therefore 1 set of 5 different newspapers is read by 60 students.
So total 5 sets are required to make 300 students read all 5 newspapers.
Therefore 5sets=5×5=25 newspapers
 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

total different newspapers possible 300 x 5 = 1500 newspapers. total different newspapers if a newspaper is read by 60 student, 1500 / 60 = 25

Thanks and Regards